OpenStudy (anonymous):

2 1 _____ + _____ = 6/x+5 x + 2 5

OpenStudy (anonymous):

x = 13?

jimthompson5910 (jim_thompson5910):

If the equation is \[\Large \frac{2}{x+2} + \frac{1}{5} = \frac{6}{x+5}\] x = 13 is just one of the two solutions

OpenStudy (anonymous):

x= 0 & x= 13

jimthompson5910 (jim_thompson5910):

better, x = 0 is also a solution

OpenStudy (anonymous):

youre confusing me

OpenStudy (anonymous):

was i right?

jimthompson5910 (jim_thompson5910):

x = 0 or x = 13 are both solutions

OpenStudy (anonymous):

dopeeee! xoxoxoxoo

jimthompson5910 (jim_thompson5910):

either one works if you replace each x with that number

jimthompson5910 (jim_thompson5910):

make sense?

OpenStudy (anonymous):

yes, it actually does.

jimthompson5910 (jim_thompson5910):

ok just checking